[figure: 9.7 cm]

NEGATIVE MAGNETORESISTANCE OF LAYERED HTSC COMPOUNDS

Yu.V. Medvedev^{a}, S.V. Yampolskii^{a,b} and S.J. Lewandowski^{c} a A.A.Galkin Donetsk Physics & Technology Institute of the Ukrainian National

Academy of Sciences, Donetsk 340114, Ukraine

b Department of Physics, University of Antwerp (UIA), B-2610 Antwerpen, Belgium c Instytut Fizyki PAN, Al.Lotnikow,32, PLO2-668 Warszawa, Poland

Magnetoresistance (MR) of layered high-T_{c} systems with quasi-two- dimensional superconductivity in magnetic field applied parallel to superconducting (S) planes has been investigated theoretically. In such configuration an increase of the magnetic field H value decreases the critical temperature of S regions [1]. It causes both a suppression of S gap parameter and an increase of tunnel current between the S planes and, consequently, leads to negative MR. MR dependence on H has been calculated for different values of ratio between S layer thickness d_{S} and the normal layer thickness d_{N}. It is found that in the d_{S} ∼^{<} d_{N} case there exists a region on the H-T phase diagram where MR is positive. In the opposite case (d_{S} ≫ d_{N}) MR is negative anywhere on the H-T diagram. These results can be used also for analysis of current-voltage characteristics of both S-I-S and S-I-N tunnel junctions.

[1] V.N.Varyukhin, G.G.Levchenko, Yu.V.Medvedev, A.V.Pashchenko, and S.J.Lewandowski, Low Temp. Phys. 24, 234 (1998).

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : S.V.Yampolskii

Address for correspondence : Department of Physics University of Antwerp(UIA) Universiteitsplein 1 B-2610 Antwerpen, Belgium

Email address : yampolsk@uia.ua.ac.be